Listed below are paired data consisting of amounts spent on advertising (in millions of dollars) and the profits (in millions of dollars). Determine if there is a significant positive linear correlation between advertising cost and profit . Use a significance level of 0.05 and round all values to 4 decimal places.

Advertising Cost Profit
3 14
4 17
5 20
6 27
7 17
8 31
9 27

Ho: ρ = 0
Ha: ρ > 0

Find the Linear Correlation Coefficient
r =

Find the p-value
p-value =   

The p-value is

  • Greater than αα
  • Less than (or equal to) αα

The p-value leads to a decision to

  • Do Not Reject Ho
  • Reject Ho
  • Accept Ho

The conclusion is

  • There is a significant negative linear correlation between advertising expense and profit.
  • There is insufficient evidence to make a conclusion about the linear correlation between advertising expense and profit.
  • There is a significant positive linear correlation between advertising expense and profit.
  • There is a significant linear correlation between advertising expense and profit.
Answer #1
X Y XY
3 14 42 9 196
4 17 68 16 289
5 20 100 25 400
6 27 162 36 729
7 17 119 49 289
8 31 248 64 961
9 27 243 81 729
Ʃx = 42
Ʃy = 153
Ʃxy = 982
Ʃx² = 280
Ʃy² = 3593
Sample size, n = 7
x̅ = Ʃx/n = 42/7 = 6
y̅ = Ʃy/n = 153/7 = 21.85714286
SSxx = Ʃx² - (Ʃx)²/n = 280 - (42)²/7 = 28
SSyy = Ʃy² - (Ʃy)²/n = 3593 - (153)²/7 = 248.8571429
SSxy = Ʃxy - (Ʃx)(Ʃy)/n = 982 - (42)(153)/7 = 64

Null and alternative hypothesis:

Ho: ρ = 0

Ha: ρ > 0

Correlation coefficient, r = SSxy/√(SSxx*SSyy) = 64/√(28*248.85714) = 0.7667

Test statistic :  

t = r*√(n-2)/√(1-r²) = 0.7667 *√(7 - 2)/√(1 - 0.7667²) = 2.6704

df = n-2 = 5

p-value = T.DIST.RT(2.6704, 5) = 0.0222

Conclusion:

p-value < α Reject the null hypothesis.

Conclusion :

  • There is a significant positive linear correlation between advertising expense and profit.